Find the equation of the straight line with gradient -1/2 that passes through the point (5,3)


1
Expert's answer
2022-02-06T13:17:42-0500

The formula for equation of a line given the Slope "m" and point "(x_1,y_1)" is;

"\displaystyle
m=\frac{y-y_1}{x-x_1}"

Here, "\displaystyle
m=-\frac{1}{2},\text{ and}(x_1,y_1)=(5,3)." Thus we have the equation of the line as;

"\displaystyle
-\frac{1}{2}=\frac{y-3}{x-5}\Rightarrow y=-\frac{x}{2}+\frac{11}{2}"